# s1: Simple test-time scaling

### v2

@@ 2.1 Initial collection of 59K samples
We collect an initial 59,029 questions from 16 diverse sources following three guiding principles. Quality: Datasets should be of high quality; we always inspect samples and ignore datasets with, e.g., poor formatting; Difficulty: Datasets should be challenging and require significant reasoning effort; Diversity: Datasets should stem from different fields to cover different reasoning tasks. We collect datasets of two categories:
@@
Our largest source is NuminaMATH (LI et al., 2024) with 30,660 mathematical problems from online websites. We also include historical AIME problems (1983-2021). To enhance diversity, we add OlympicArena (Huang et al., 2024a) with 4,250 questions spanning Astronomy, Biology, Chemistry, Computer Science, Geography, Mathematics, and Physics from various Olympiads. OmniMath (Gao et al., 2024a) adds 4,238 competition-level mathematics problems. We also include 2,385 problems from AGIEval (Zhong et al., 2023), which features questions from standardized tests like SAT and LSAT, covering English, Law, and Logic. We refer to Table 6 in §B for our other sources.
@@
For each question, we generate a reasoning trace and solution using the Google Gemini Flash Thinking API (Google, 2024) extracting its reasoning trace and response. This yields 59K triplets of a question, generated reasoning trace, and generated solution. Examples from our dataset are in §C.2. We decontaminate all samples against our evaluation questions (MATH500, GPQA Diamond, AIME24; §B.5) using 8-grams and deduplicate the data.
@@ 2.2 Final selection of 1K samples
We first remove any questions where we ran into any API errors reducing our dataset to 54,116 samples. Next, we filter out low-quality examples by checking if they contain any string patterns with formatting issues, such as ASCII art diagrams, non-existent image references, or inconsistent question numbering reducing our dataset to 51,581 examples. From this pool, we identify 384 samples for our final 1,000 samples from datasets that we perceive as high-quality and not in need of further filtering (see §B.4 for details).
@@
For difficulty, we use two indicators: model performance and reasoning trace length. We evaluate two models on each question: Qwen2.5-7B-Instruct and Qwen2.5-32B-Instruct (Qwen et al., 2024), with correctness assessed by Claude 3.5 Sonnet comparing each attempt against the reference solution (see §B.3 for the grading protocol). We measure the token length of each reasoning trace to indicate problem difficulty using the Qwen2.5 tokenizer. This relies on the assumption that more difficult problems require more thinking tokens. Based on the grading, we remove questions that either Qwen2.5-7B-Instruct or Qwen2.5-32B-Instruct can solve correctly and thus may be too easy. By using two models we reduce the likelihood of an easy sample slipping through our filtering due to a rare mistake on an easy question of one of the models. This brings our total samples down to 24,496, setting the stage for the next round of subsampling based on diversity. While filtering with these two models may be optimized for our setup as we will also use Qwen2.5-32B-Instruct as our model to finetune, the idea of model-based filtering generalizes to other setups.
@@
To quantify diversity we classify each question into specific domains using Claude 3.5 Sonnet based on the Mathematics Subject Classification (MSC) system (e.g., geometry, dynamic systems, real analysis, etc.) from the American Mathematical Society. [^1] https://mathscinet.ams.org/mathscinet/msc/msc2020.html The taxonomy focuses on topics in mathematics but also includes other sciences such as biology, physics, and economics. To select our final examples from the pool of 24,496 questions, we first choose one domain uniformly at random. Then, we sample one problem from this domain according to a distribution that favors longer reasoning traces (see §B.4 for details) as motivated in Difficulty. We repeat this process until we have 1,000 total samples.
This three-stage process yields a dataset spanning 50 different domains (see Table 5). In §5.1, we will show that using our three criteria in combination is important, as only relying on quality, diversity, or difficulty in isolation leads to worse datasets. Examples from our dataset are in §C.2.
@@ 3.1 Method
We propose a simple decoding-time intervention by forcing a maximum and/or minimum number of thinking tokens at test time. Specifically, we enforce a maximum token count by simply appending the end-of-thinking token delimiter and “Final Answer:” to early exit the thinking stage and make the model provide its current best answer. To enforce a minimum, we suppress the generation of the end-of-thinking token delimiter and optionally append the string “Wait” to the model’s current reasoning trace to encourage the model to reflect on its current generation. Figure 3 contains an example of how this simple approach can lead the model to arrive at a better answer.
@@
We benchmark budget forcing with: (I) Conditional length-control methods, which rely on telling the model in the prompt how long it should generate for. We group them by granularity into (a) Token-conditional control: We specify an upper bound of thinking tokens in the prompt; (b) Step-conditional control: We specify an upper bound of thinking steps, where each step is around 100 tokens; (c) Class-conditional control: We write two generic prompts that tell the model to either think for a short or long amount of time (see §D.1 for details). (II) Rejection sampling, which samples until a generation fits a predetermined compute budget. This oracle captures the posterior over responses conditioned on its length.
@@ 4.1 Setup
We perform supervised finetuning on Qwen2.5-32B-Instruct using s1K to obtain our model s1-32B using basic hyperparameters outlined in §C. Finetuning took 26 minutes on 16 NVIDIA H100 GPUs with PyTorch FSDP.
@@
We select three representative reasoning benchmarks widely used in the field: AIME24 (of America, 2024) consists of 30 problems that were used in the 2024 American Invitational Mathematics Examination (AIME) held from Wednesday, January 31 – Thursday, February 1, 2024. AIME tests mathematical problem-solving with arithmetic, algebra, counting, geometry, number theory, probability, and other secondary school math topics. High-scoring high school students in the test are invited to participate in the United States of America Mathematics Olympiad (USAMO). All AIME answers are integers ranging from $000$ to $999$, inclusive. Some AIME problems rely on figures that we provide to our model using the vector graphics language Asymptote as it cannot take image inputs. MATH500 (Hendrycks et al., 2021) is a benchmark of competition math problems of varying difficulty. We evaluate on the same 500 samples selected by OpenAI in prior work (Lightman et al., 2023). GPQA Diamond (Rein et al., 2023) consists of 198 PhD-level science questions from Biology, Chemistry and Physics. Experts with PhDs in the corresponding domains only achieved 69.7% on GPQA Diamond (OpenAI, 2024). When we write “GPQA” in the context of evaluation in this work, we always refer to the Diamond subset. We build on the “lm-evaluation-harness” framework (Gao et al., 2021; Biderman et al., 2024).
@@
We benchmark s1-32B against: OpenAI o1 series (OpenAI, 2024), which are closed-source models that popularized the idea of test-time scaling; DeepSeek r1 series (Team, 2024a), which are open-weight reasoning models with up to o1-level performance, concurrently released to ours; Qwen’s QwQ-32B-preview (Team, 2024b), a 32B open-weight reasoning model without disclosed methodology; Sky-T1-32B-Preview (Team, 2025) and Bespoke-32B (Labs, 2025), which are open models with open reasoning data distilled from QwQ-32B-preview and r1; Google Gemini 2.0 Flash Thinking Experimental (Google, 2024), the API that we distill from. As it has no official evaluation scores, we use the Gemini API to benchmark it ourselves. However, the “recitation error” of the Gemini API makes evaluation challenging. [^2] https://github.com/google/generative-ai-docs/issues/257 We circumvent this, by manually inserting all 30 AIME24 questions in its web interface where the error does not appear. However, we leave out MATH500 (500 questions) and GPQA Diamond (198 questions), thus they are N.A. in LABEL:tab:perf. Our model, s1-32B, is fully open including weights, reasoning data, and code.
Table 1: s1-32B is an open and sample-efficient reasoning model. We evaluate s1-32B, Qwen, and Gemini (some entries are unknown (N.A.), see §4). Other results are from the respective reports (Qwen et al., 2024; Team, 2024b; OpenAI, 2024; DeepSeek-AI et al., 2025; Labs, 2025; Team, 2025). # ex. = number examples used for reasoning finetuning; BF = budget forcing.
@@ 4.2 Performance
In Figure 2 (right) and LABEL:tab:perf we compare s1-32B with other models. We find that s1-32B is the most sample-efficient open data reasoning model. It performs significantly better than our base model (Qwen2.5-32B-Instruct) despite just training it on an additional 1,000 samples. The concurrently released r1-32B shows stronger performance than s1-32B while also only using SFT (DeepSeek-AI et al., 2025). However, it is trained on 800 $\times$ more reasoning samples. It is an open question whether one can achieve their performance with just 1,000 samples. Finally, our model nearly matches Gemini 2.0 Thinking on AIME24. As s1-32B is distilled from Gemini 2.0, this shows our distillation procedure was likely effective.
@@ 5.2 Test-time scaling methods
Class-conditional control We provide benchmark scores for this method in §D.1 and summarize three findings here: (1) Token-conditional control fails without budget forcing, as our model cannot reliably count tokens - even when trained to do so. (2) Under step-conditional control, the model generates a similar total number of tokens when given different step targets, as the model goes from few steps with many tokens per step, to many steps with few tokens in each step. Thus, the model learns to hack its way around the compute constraint making the controllability of this method mediocre. (3) Class-conditional control can work - telling a model to simply think longer can increase its test-time compute and performance, which leads good scaling in Table 3.
@@
Surprisingly, we find that simply sampling until the generation fits a specific length leads to an inverse scaling trend as depicted in Figure 6. In §D.2 we inspect a question, which was answered correctly by the model when rejection sampling for $\leq 4000$, but not for the $\leq 8000$ token setting. In the $\leq 4000$ setting the model directly jumps to the correct approach, while for the $\leq 8000$ setting it backtracks a lot. We hypothesize that there is a correlation such that shorter generations tend to be the ones where the model was on the right track from the start, whereas longer ones tend to be ones where the model made mistakes and thus backtracks or questions itself. This leads to longer samples often being wrong when rejection sampling and thus the inverse scaling trend.
@@ 6.1 Sample-efficient reasoning
To evaluate and push the limits of these models, increasingly challenging benchmarks have been introduced, such as Olympiad-level science competitions (He et al., 2024; Jain et al., 2024; Zhong et al., 2023) and others (Srivastava et al., 2023; Glazer et al., 2024; Su et al., 2024; Kim et al., 2024; Phan et al., 2025). To enhance models’ performance on reasoning-related tasks, researchers have pursued several strategies: Prior works have explored continuing training language models on specialized corpora related to mathematics and science (Azerbayev et al., 2023; Yang et al., 2024), sometimes even synthetically generated data (Yu et al., 2024). Others have developed training methodologies specifically aimed at reasoning performance (Zelikman et al., 2022, 2024; Luo et al., 2025; Yuan et al., 2025; Wu et al., 2024a). Another significant line of work focuses on prompting-based methods to elicit and improve reasoning abilities, including methods like Chain-of-Thought prompting (Wei et al., 2023; Yao et al., 2023a, b; Bi et al., 2023; Fu et al., 2023; Zhang et al., 2024b; Xiang et al., 2025; Hu et al., 2024). These combined efforts aim to advance the reasoning ability of language models, enabling them to handle more complex and abstract tasks effectively.
@@ Methods
As we introduce in §3, we differentiate two methods to scale test-time compute: parallel and sequential. The former relies on multiple solution attempts generated in parallel and selecting the best outcome via specific criteria. These criteria include choosing the most frequent response for majority voting or the best response based on an external reward for Best-of-N (Brown et al., 2024; Irvine et al., 2023; Snell et al., 2024). Unlike repeated sampling, previous sequential scaling methods let the model generate solution attempts sequentially based on previous attempts, allowing it to refine each attempt based on previous outcomes (Snell et al., 2024; Hou et al., 2025; Lee et al., 2025). Tree-based search methods (Gandhi et al., 2024; Wu et al., 2024b) offer a hybrid approach between sequential and parallel scaling, such as Monte-Carlo Tree Search (MCTS) (Liu et al., 2024; Zhang et al., 2023; Zhou et al., 2024; Choi et al., 2023) and guided beam search (Xie et al., 2023). REBASE (Wu et al., 2024b) employs a process reward model to balance exploitation and pruning during tree search. Empirically, REBASE has been shown to outperform sampling-based methods and MCTS (Wu et al., 2024b). Reward models (Lightman et al., 2023; Wang et al., 2024b, c) play a key role in these methods. They come in two variants: outcome reward models and process reward models. Outcome reward models (Xin et al., 2024; Ankner et al., 2024) assign a score to complete solutions and are particularly useful in Best-of-N selection, while process reward models (Lightman et al., 2023; Wang et al., 2024b; Wu et al., 2024b) assess individual reasoning steps and are effective in guiding tree-based search methods.

### v3

@@ 2.1 Initial collection of 59K samples
We collect an initial 59,029 questions from 16 sources following three guiding principles. Quality: Datasets should be high-quality; we always inspect samples and ignore datasets with, e.g., poor formatting; Difficulty: Datasets should be challenging and require significant reasoning effort; Diversity: Datasets should stem from various fields to cover different reasoning tasks. We collect datasets of two categories:
@@
Our largest source is NuminaMATH (LI et al., 2024) with 30,660 mathematical problems from online websites. We also include historical AIME problems (1983-2021). To enhance diversity, we add OlympicArena (Huang et al., 2024a) with 4,250 questions spanning Astronomy, Biology, Chemistry, Computer Science, Geography, Mathematics, and Physics from various Olympiads. OmniMath (Gao et al., 2024a) adds 4,238 competition-level mathematics problems. We also include 2,385 problems from AGIEval (Zhong et al., 2023), which features questions from standardized tests like SAT and LSAT, covering English, Law, and Logic. We refer to Table 7 in §C for our other sources.
@@
For each question, we generate a reasoning trace and solution using the Google Gemini Flash Thinking API (Google, 2024) extracting its reasoning trace and response. This yields 59K triplets of a question, generated reasoning trace, and generated solution. Examples from our dataset are in §D.2. We decontaminate all samples against our evaluation questions (MATH500, GPQA Diamond, AIME24; §C.5) using 8-grams and deduplicate the data.
@@ 2.2 Final selection of 1K samples
We first remove any questions where we ran into any API errors reducing our dataset to 54,116 samples. Next, we filter out low-quality examples by checking if they contain any string patterns with formatting issues, such as ASCII art diagrams, non-existent image references, or inconsistent question numbering reducing our dataset to 51,581 examples. From this pool, we identify 384 samples for our final 1,000 samples from datasets that we perceive as high-quality and not in need of further filtering (see §C.4 for details).
@@
For difficulty, we use two indicators: model performance and reasoning trace length. We evaluate two models on each question: Qwen2.5-7B-Instruct and Qwen2.5-32B-Instruct (Qwen et al., 2024), with correctness assessed by Claude 3.5 Sonnet comparing each attempt against the reference solution (see §C.3 for the grading protocol). We measure the token length of each reasoning trace to indicate problem difficulty using the Qwen2.5 tokenizer. This relies on the assumption that more difficult problems require more thinking tokens. Based on the grading, we remove questions that either Qwen2.5-7B-Instruct or Qwen2.5-32B-Instruct can solve correctly and thus may be too easy. By using two models we reduce the likelihood of an easy sample slipping through our filtering due to a rare mistake on an easy question of one of the models. This brings our total samples down to 24,496, setting the stage for the next round of subsampling based on diversity. While filtering with these two models may be optimized for our setup as we will also use Qwen2.5-32B-Instruct as our model to finetune, the idea of model-based filtering generalizes to other setups.
@@
To quantify diversity, we classify questions into domains using Claude 3.5 Sonnet based on the Mathematics Subject Classification (MSC) system (e.g., geometry, combinatorics, etc.) from the American Mathematical Society. [^1] https://mathscinet.ams.org/mathscinet/msc/msc2020.html The taxonomy focuses on topics in mathematics but also includes other sciences such as biology, physics, and economics. To select our final examples from the pool of 24,496 questions, we first choose one domain uniformly at random. Then, we sample one problem from this domain according to a distribution that favors longer reasoning traces (see §C.4 for details) as motivated in Difficulty. We repeat this process until we have 1,000 total samples spanning 50 domains.
In §5.1, we will show that using our three criteria in combination is important, as only relying on quality, diversity, or difficulty in isolation leads to worse datasets. Some distilled generations are incorrect, which we allow in our data as we focus on capturing the reasoning process rather than entirely correct solutions. Our grader (§C.3) deems 53.6% correct in s1K and 63.0% in our follow-up s1K-1.1 (see §A).
@@ 3.1 Method
We propose a simple decoding-time intervention by forcing a maximum and/or minimum number of thinking tokens. Specifically, we enforce a maximum token count by simply appending the end-of-thinking token delimiter and optionally “Final Answer:” to early exit the thinking stage and make the model provide its current best answer. To enforce a minimum, we suppress the generation of the end-of-thinking token delimiter and optionally append the string “Wait” to the model’s current reasoning trace to encourage the model to reflect on its current generation. Figure 3 contains an example of how this simple approach can lead the model to arrive at a better answer.
@@
We benchmark budget forcing with: (I) Conditional length-control methods, which rely on telling the model in the prompt how long it should generate for. We group them by granularity into (a) Token-conditional control: We specify an upper bound of thinking tokens in the prompt; (b) Step-conditional control: We specify an upper bound of thinking steps, where each step is around 100 tokens; (c) Class-conditional control: We write two generic prompts that tell the model to either think for a short or long amount of time (see §E.1 for details). (II) Rejection sampling, which samples until a generation fits a predetermined compute budget. This oracle captures the posterior over responses conditioned on its length.
@@ 4.1 Setup
We perform supervised finetuning on Qwen2.5-32B-Instruct using s1K to obtain our model s1-32B using basic hyperparameters outlined in §D. Finetuning took 26 minutes on 16 NVIDIA H100 GPUs with PyTorch FSDP.
@@
We select three representative reasoning benchmarks widely used in the field: AIME24 (of America, 2024) has 30 problems that were used in the 2024 American Invitational Mathematics Examination (AIME) held from January 31 – February 1, 2024. AIME tests mathematical problem-solving with arithmetic, algebra, counting, geometry, number theory, probability, and other secondary school math topics. High-scoring high school students in the test are invited to participate in the United States of America Mathematics Olympiad (USAMO). All AIME answers are integers ranging from $000$ to $999$, inclusive. Some AIME problems rely on figures that we provide to our model using the vector graphics language Asymptote as it cannot take image inputs. MATH500 (Hendrycks et al., 2021) is a benchmark of competition math problems of varying difficulty. We evaluate on the same 500 samples selected by OpenAI in prior work (Lightman et al., 2023). GPQA Diamond (Rein et al., 2023) consists of 198 PhD-level science questions from Biology, Chemistry and Physics. Experts with PhDs in the corresponding domains only achieved 69.7% on GPQA Diamond (OpenAI, 2024). When we write “GPQA” in the context of evaluation in this work, we always refer to the Diamond subset. We build on the “lm-evaluation-harness” framework (Gao et al., 2021; Biderman et al., 2024). Unless otherwise specified, we evaluate with a temperature of 0 (greedy) and measure accuracy (equivalent to pass@1).
@@
We benchmark s1-32B against: OpenAI o1 series (OpenAI, 2024), closed-source models that popularized test-time scaling; DeepSeek r1 series (DeepSeek-AI et al., 2025), open-weight reasoning models with up to o1-level performance; Qwen’s QwQ-32B-preview (Team, 2024), a 32B open-weight reasoning model without disclosed methodology; Sky-T1-32B-Preview (Team, 2025) and Bespoke-32B (Labs, 2025), open models with open reasoning data distilled from QwQ-32B-preview and r1; Google Gemini 2.0 Flash Thinking Experimental (Google, 2024), the API that we distill from. As it has no official evaluation scores, we use the Gemini API to benchmark it ourselves. However, the “recitation error” of the Gemini API makes evaluation challenging. [^2] https://github.com/google/generative-ai-docs/issues/257 We circumvent this, by manually inserting all 30 AIME24 questions in its web interface where the error does not appear. However, we leave out MATH500 (500 questions) and GPQA Diamond (198 questions), thus they are N.A. in LABEL:tab:perf. Our model, s1-32B, is fully open including weights, reasoning data, and code.
Table 1: s1-32B is a strong open reasoning model. We evaluate s1-32B, Qwen, and Gemini (some entries are unknown (N.A.), see §4). Other results are from the respective reports (Qwen et al., 2024; Team, 2024; OpenAI, 2024; DeepSeek-AI et al., 2025; Labs, 2025; Team, 2025). # ex. = number examples used for reasoning finetuning; BF = budget forcing. See §A for our better s1.1 model.
@@ 4.2 Performance
In Figure 2 (right) and LABEL:tab:perf we compare s1-32B with other models. We find that s1-32B is the most sample-efficient open data reasoning model. It performs significantly better than our base model (Qwen2.5-32B-Instruct) despite just training it on an additional 1,000 samples. The concurrently released r1-32B shows stronger performance than s1-32B while also only using SFT (DeepSeek-AI et al., 2025). However, it is trained on 800 $\times$ more reasoning samples. It is an open question whether one can achieve their performance with just 1,000 samples. Finally, our model nearly matches Gemini 2.0 Thinking on AIME24. As the data for s1-32B is distilled from Gemini 2.0, this shows our distillation procedure was likely effective.
@@ 5.2 Test-time scaling methods
Class-conditional control We provide benchmark scores for this method in §E.1 and summarize three findings here: (1) Token-conditional control fails without budget forcing, as our model cannot reliably count tokens - even when trained to do so. (2) Under step-conditional control, the model generates a similar total number of tokens when given different step targets, as the model goes from few steps with many tokens per step, to many steps with few tokens in each step. Thus, the model learns to hack its way around the compute constraint making the controllability of this method mediocre. (3) Class-conditional control can work - telling a model to simply think longer can increase its test-time compute and performance, which leads good scaling in Table 3.
@@
Surprisingly, we find that simply sampling until the generation fits a specific length leads to an inverse scaling trend as depicted in Figure 6. In §E.2 we inspect a question, which was answered correctly by the model when rejection sampling for $\leq 4000$, but not for the $\leq 8000$ token setting. In the $\leq 4000$ setting the model directly jumps to the correct approach, while for the $\leq 8000$ setting it backtracks a lot. We hypothesize that there is a correlation such that shorter generations tend to be the ones where the model was on the right track from the start, whereas longer ones tend to be ones where the model made mistakes and thus backtracks or questions itself. This leads to longer samples often being wrong when rejection sampling and thus the inverse scaling trend.
@@ 6.1 Sample-efficient reasoning
To evaluate and push the limits of these models, increasingly challenging benchmarks have been introduced, such as Olympiad-level science competitions (He et al., 2024; Jain et al., 2024; Zhong et al., 2023) and others (Srivastava et al., 2023; Glazer et al., 2024; Su et al., 2024; Kim et al., 2024; Phan et al., 2025). To enhance models’ performance on reasoning-related tasks, researchers have pursued several strategies: Prior works have explored continuing training language models on specialized corpora related to mathematics and science (Azerbayev et al., 2023; Yang et al., 2024), sometimes even synthetically generated data (Yu et al., 2024). Others have developed training methodologies specifically aimed at reasoning performance (Zelikman et al., 2022, 2024; Luo et al., 2025; Yuan et al., 2025; Wu et al., 2024a). Another significant line of work focuses on prompting-based methods to elicit and improve reasoning abilities, including methods like Chain-of-Thought prompting (Wei et al., 2023; Yao et al., 2023a, b; Bi et al., 2023; Fu et al., 2023; Zhang et al., 2024b; Xiang et al., 2025; Hu et al., 2024; Diao et al., 2024). These combined efforts aim to advance the reasoning ability of language models, enabling them to handle more complex and abstract tasks effectively.
@@ Methods
As we introduce in §3, we differentiate two methods to scale test-time compute: parallel and sequential. The former relies on multiple solution attempts generated in parallel and selecting the best outcome via specific criteria. These criteria include choosing the most frequent response for majority voting or the best response based on an external reward for Best-of-N (Brown et al., 2024; Irvine et al., 2023; Levi, 2024). Unlike repeated sampling, previous sequential scaling methods let the model generate solution attempts sequentially based on previous attempts, allowing it to refine each attempt based on previous outcomes (Snell et al., 2024; Hou et al., 2025; Lee et al., 2025). Tree-based search methods (Gandhi et al., 2024; Wu et al., 2024b) offer a hybrid approach between sequential and parallel scaling, such as Monte-Carlo Tree Search (MCTS) (Liu et al., 2024; Zhang et al., 2023; Zhou et al., 2024; Choi et al., 2023) and guided beam search (Xie et al., 2023). REBASE (Wu et al., 2024b) employs a process reward model to balance exploitation and pruning during tree search. Empirically, REBASE has been shown to outperform sampling-based methods and MCTS (Wu et al., 2024b). Reward models (Lightman et al., 2023; Wang et al., 2024b, c) play a key role in these methods. They come in two variants: outcome reward models and process reward models. Outcome reward models (Xin et al., 2024; Ankner et al., 2024) assign a score to complete solutions and are particularly useful in Best-of-N selection, while process reward models (Lightman et al., 2023; Wang et al., 2024b; Wu et al., 2024b) assess individual reasoning steps and are effective in guiding tree-based search methods.